# Beyond the Carnot limit: work extraction via an entropy battery

### v1

@@ II.1 Specifics of the entropy battery
The specific design of the entropy battery is shown in Fig.2. The battery contains two overlapping ensembles, one small ensemble that acts as the working fluid (WF) for the Raman heat engine, and the other being a much larger reservoir (RES) of spin-polarized, energetically cold particles.
@@
The working fluid consists of $N_{\text{wf}}$ spin-1 particles, and the reservoir contains $N_{\text{res}}$ spin-$S$ particles. These ensembles continuously thermalize in both spin and energy degrees of freedom via spin and energy exchange collisions, such that at any moment in time $T_{\text{wf}}\approx T_{\text{res}}$ and $\tau_{\text{wf}}\approx\tau_{\text{res}}$. Since $N_{\text{res}}\gg N_{\text{wf}}$, we approximate the total entropy of the entropy battery as arising primarily from the reservoir, $\mathbb{S}_{\text{battery}}=\mathbb{S}_{\text{wf}}+\mathbb{S}_{\text{res}}\approx\mathbb{S}_{\text{res}}$. The entropy capacity of the battery depends on the number of spin and energy states, as well as the number of particles, as dictated by eqs.(2)–(2). We now introduce the environment ensemble, which possesses only an energy degree of freedom. The joint environment, entropy battery system is assumed to be isolated from any ‘grander’ system, such that the composite system is closed, and there is a net zero change in entropy during the environment-battery thermalization process. As the environment couples to the entropy battery only via the energy degree of freedom, then the total change in entropy must be constrained by
$\Delta\mathbb{S}_{\text{env}}(\tau_{\text{env}},\tau_{f})+\int_{\tau_{E,0}}^{\tau_{f}}\frac{1}{\tau_{E}}C_{V}d\tau_{E}+\int_{\tau_{s,0}}^{\tau_{f}}\frac{1}{\tau_{s}}C_{s}d\tau_{s}=0,$ (24)
where $\tau_{E,0}$ and $\tau_{s,0}$ are the battery’s unitless energy and spin initial temperatures, and $\tau_{\text{env}}$ is the environments initial temperature, where $\tau_{\text{env}}>\tau_{E,0},\tau_{s,0}$. The unitless energetic temperature is $\tau_{E}=Tk_{B}/h\nu$, see Appendix E.3. $\tau_{f}$ is the unitless final temperature once equilibrium is reached, and is common across the conserved quantity waste baths as required by the zero’th law (2). The final temperature is dependent on the entropic responses $C_{s}/\tau$ of the environment, and the energy, spin waste baths of the battery.
II.2 Results
As we are interested in the theoretical potential of work extraction using the entropy battery, then we should use particle statistics that maximize the battery’s entropy capacity. At face value, it would appear as though distinguishable particles would provide the highest entropy capacity according to eq.(2), however, as will be discussed in §III, distinguishability implies non-interaction, making distinguishable particles unrealistic as our model (and real systems) require interactions between the canonical ensembles for entropy to transfer. The next best case is an ensemble of bosons. If both the environment and the battery have a large number of bosons, on the order of Avogadro’s number or more, then the waste responses $C_{s},C_{V}$ in the integrals of eq.(24) are approximately eq.(2), and the indefinite integrals $\int C_{s}/\tau d\tau$ evaluate to
@@ II.2 Results
We use this expression to numerically solve for the final temperature by minimising $|\Delta\mathbb{S}|$ in eq.(24).
The final temperature is used in the analytic boson heat expression (2) to calculate the energetic heat transferred from the environment to the energy bath of the battery, and therefore the energetic work liberated. This represents the unitless work that is made available during a conventional heat engine cycle
$W_{\text{conventional}}=|Q_{\text{env}}|-|Q_{\text{batt}}|.$
The spin bath also absorbs entropy via the coupling heat engine, while simultaneously liberating heat energy in the battery as energetic work, at the cost of spin labor [10, 11]. Therefore, the total unitless energetic work of the entropy battery becomes
$W_{\text{battery}}=|Q_{\text{env}}|-|Q_{\text{batt}}|+|\mathcal{Q}_{\text{batt}}|$ (26)
where $\mathcal{Q}_{\text{batt}}$ is the spin therm absorbed by the battery’s spin waste bath. These heats and spin equivalents are all dependent on their respective waste responses $C_{s}$, and so the number of states each waste bath has access to. The energy baths of both the environment and the battery have in principal an infinite number of states as they are thermal motional states. We use a truncated energy state space ($d_{E}=400$) in our analysis without impacting accuracy.
@@
$\eta_{\text{energy}}=\frac{W_{\text{battery}}}{|Q_{\text{env}}|}=1-\frac{|Q_{\text{batt}}|-|\mathcal{Q}_{\text{batt}}|}{|Q_{\text{env}}|},$ (27)
@@
and the results are shown in Fig.3 for varying initial battery temperatures. Fig.3 shows that once $d_{s}\geq 2$, the energy efficiency surpasses the Carnot limit. At 5 spin states, all initial temperatures considered reach close to $100\%$ energy efficiency. This occurs despite all waste baths being initially thermal states, without induced quantum coherence. Further, as the entropy transfer occurs adiabatically, the thermalization process must be endoreversible. The energy efficiency with 0 spin states follows $1-\sqrt{\tau_{1}/\tau_{2}}$, meaning our model represents the expected efficiency one would obtain when running a heat engine at maximum power [39, 40, 41]. See Appendix G.
@@
Figure 3: Efficiency of energetic work extraction $\eta_{\text{energy}}$ for varying number of spin states per particle calculated using the entropy and heat equations (25) and (2). In these simulations we used $d_{\text{env}}=d_{\text{ens},E}=400$ without loss in accuracy. Also shown are 4 sets of initial temperatures for the ensembles energy and spin baths, with $\tau_{\text{batt}}=\tau_{E,0}=\tau_{s,0}$, while the environments initial temperature was set to $\tau_{\text{env}}=0.6$ for cases. These unitless temperatures were chosen as it would correspond to a temperature range between room temperature (300 K) for $\tau_{\text{batt}}=0.3$ on the low end and 600 K $\tau_{\text{env}}=0.6$ on the high end with a Debye cut off frequency of 2121 THz, which is a typical maximum frequency for metals [37]. These temperature ranges are typical in coal and nuclear power plants, and concentrated solar farms. As the number of spin states increases, the energy efficiency approaches 100%, with all initial conditions reaching approximately this limit at 5 states. This is well above the Carnot efficiency limit $\eta_{\text{Carnot}}=1-\tau_{\text{batt}}/\tau_{\text{env}}$ which is shown by the red bars when $d=0$, with values between 16%-50% for the varying initial ensemble temperatures.
II.3 Discussion
@@
The Carnot limit sets the maximum efficiency achievable by a heat engine operating between two thermal reservoirs of different temperatures. As mentioned in the introduction, a common strategy in quantum heat engines to surpass this limit is to induce quantum coherence within the thermal bath, either via squeezed thermal states [26, 27, 28] or other induced coherences [29, 30, 31]. All of these devices operate below the generalized Carnot limit [32], therefore there is no real quantum advantage. In contrast, the entropy battery achieves super-Carnot efficiencies in large ensembles using only maximum entropy states, characterised by their unitless temperatures, demonstrating that this limit can be surpassed only when an energetically hot source is coupled to an ensemble with multiple conserved quantities via multiple unitary heat engines.
@@
$\mathbb{W}=|Q_{\text{env}}|-|Q_{\text{batt}}|+|\mathcal{Q}_{\text{batt}}|-|\mathcal{L}|,$
where $|\mathcal{L}|=|\mathcal{Q}_{\text{batt}}|$. The generalized efficiency of the battery when considering all resource costs therefore returns to the Carnot efficiency limit $\eta=1-|Q_{\text{batt}}|/|Q_{\text{env}}|$. Eq.(27) is therefore the energetic efficiency of the battery. If the initial preparation of spin labor within the battery requires energy, and if this energy cost is equal to the energy saved during the entropy transfer process in the unitary heat engine, then the entropy battery would not have an advantageous net energy efficiency. However, charging spin-polarized reservoirs does not necessarily require an energy input. Naturally occurring sources of spin coherence can instead be used to re-polarize the reservoir, analogously to how we currently use natural cold energetic baths to cool hot ensembles (from an entropic perspective). Consequently, the entropy battery now has an additional advantage: it extends the range of usable natural resources for energetic work extraction beyond energy alone. Further, the unitary heat engine that couples the external energy and internal spin degrees of freedom is controllable in the sense that it can be switched on/off. Combining this controllability with the assumed mutual independence of the spin degree of freedom, the state of ‘charge’ of the battery would remain approximately constant even in an energetically hot environment. This makes the entropy battery useful in applications that require such controllability, with it effectively performing the job of a conventional battery, all the while storing not energy, but information in the form of coherence when charged. Such a device could find broad applications, improving the energy efficiency of conventional batteries, power plants, and vehicles, while also enabling new methods of heat storage and refrigeration. Since the battery can, in principle, be charged using naturally occurring sources of quantum coherence in arbitrary conserved quantities, it opens the way to a future in which information itself becomes the only resource, a resource that can be controllably exchanged between conserved quantities as desired.
@@
This paper focuses on an ideal scenario in which the battery-environment systems operate adiabatically. We briefly note the role of open-system effects here. Irreversibility may arise from the injection of entropy from an external source during thermalization, or from non-independent conserved quantities, such that the total entropy is non-extensive, $\mathbb{S}_{\text{battery}}<\mathbb{S}_{\text{energy}}+\mathbb{S}_{\text{spin}}$. Both reduce the effective entropy capacity of the battery, increasing the final temperature once energy and spin equilibrate across the conserved quantities and ensembles, and decreases the extractable work. Beyond improving isolation or enhancing the mutual independence of the conserved quantities, one could compensate for this loss by increasing the number of states, however that analysis is beyond the scope of this paper.
@@
Another particular field that may apply the concepts developed in this paper is quantum error correction. The decoherence of a qubit over time can be described as a thermalization process with a thermal environment [42, 43], and as such, one can imagine that having access to another conserved quantity waste bath that is independent of the basis of measurement (the qubit) within a single physical system (the ion, superconductor etc.) would allow for the ‘re-direction’ of decoherence into the unused degree of freedom. If possible, this would slow the decoherence time of the measured qubit, potentially leading to less dependence on error correcting qubits. Actual implementation of such an error correcting technique is beyond the scope of this paper, however NMR algorithmic cooling [20, 19] bears some similarities in terms of using ancillary qubits (the spin states) for reversibly cooling the measurement qubit (the energy states).
III Maximizing entropy capacity
@@
The analysis presented in this paper focuses on a single operating cycle, after which the environment is left cooler than its initial state, $\tau_{f}<\tau_{\text{env}}$. Repeated operation of the battery, assuming the environment is reset to its original temperature after each cycle, would allow continuous extraction of energetic work until the battery reaches the same dimensionless temperature as the now effectively isothermal environment, $\tau_{f}=\tau_{\text{env}}$. These extra cycles increase the amount of work extractable, where the number of cycles depends on the entropy capacity of the ensemble. Therefore, it is of interest to understand how one could maximize the battery’s entropy capacity. According to eqs.(2) and (2), one could increase the number of particles NN or states $d$, as either of these increases the number of microstates $\Omega$ and thus the entropy capacity. Another way would be to increase the number of independent conserved quantities. The dependence of the entropy capacity on these variables can be understood directly by application of the maximum entropy principle [6].
@@
The probability distribution that maximizes entropy, under the constraints of normalization $\sum_{\Omega}P(\Omega)=1$, Shannon entropy (28), and the $M$ total conserved quantities
@@
where $\gamma_{m}$ and $c_{m}$ are the Lagrange multiplier and normalization constant for the $m$th conserved quantity. In this paper, all ensembles are treated as canonical, with no particle exchange and therefore no associated chemical potential. The ensemble normalization $\prod_{m}c_{m}=Z$ is,
@@
Treating the conserved quantities as mutually independent also allows let’s us focus on each conserved quantity individually. For the conserved quantity $A_{m}$, the normalized probability distribution is
@@
The microstates $\Omega_{m}$ of the $m$th conserved quantity are now independent of the other conserved quantities. Therefore, mutual independence over the conserved quantities maximizes the ensembles entropy capacity. To further increase the entropy capacity, we are limited to increasing the number of microstates associated with the $M$ conserved quantities.
III.1 Distinguishability
Figure 4: An example entropy distribution over spin temperature for the different particle statistics, distinguishable, bosonic and fermionic with 7 states. In (a) and (b) we show example distributions with $N=4$ and $N=6$ particles respectively, illustrating how distinguishable and bosonic maximum entropies increase with NN without bound, while the fermionic maximum entropy decreases as saturation is approached $N=d$ where $\mathbb{S}_{\text{Fermi}}=0$ for all temperature. Maximum entropy is achieved once $\tau=\infty$, which we have labeled for the different ensemble statistics.

### v2

@@ IV.1 Specifics of the entropy battery
The specific design of the entropy battery is shown in Fig.2. The battery contains two overlapping ensembles, one small ensemble that acts as the working fluid (WF) for the unitary heat engine (adiabatic Raman transition), and the other being a much larger reservoir (RES) of spin-polarized, energetically cold particles.
The working fluid consists of $N_{\text{wf}}$ spin-1 particles, and the reservoir contains $N_{\text{res}}$ spin-$S$ particles. These ensembles continuously thermalize in both spin and energy degrees of freedom via spin and energy exchange collisions, such that at any moment in time $T_{\text{wf}}\approx T_{\text{res}}\equiv T_{C}$ and $\tau_{\text{wf}}\approx\tau_{\text{res}}\equiv\tau_{s}$. Since $N_{\text{res}}\gg N_{\text{wf}}$, we approximate the total entropy of the entropy battery as arising primarily from the reservoir, $\mathbb{S}_{\text{battery}}=\mathbb{S}_{\text{wf}}+\mathbb{S}_{\text{res}}\approx\mathbb{S}_{\text{res}}$. Once the entropy battery couples to an energetic thermal environment with unitless temperature $\tau_{H}\equiv T_{H}k_{B}/h\nu$, as shown in Fig.1, then the joint environment-entropy battery system can begin thermalizing. The heat engine mediated coupling allows for the transfer of entropy between the energy and spin conserved quantities across the environment and battery ensembles, under the condition $\Delta\mathbb{S}=0$, therefore
$\int_{\tau_{C}}^{\tau_{f}}\frac{C_{V}}{\tau}d\tau+\int_{\tau_{s}}^{\tau_{f}}\frac{C_{s}}{\tau}d\tau=-\Delta\mathbb{S}_{\text{env}}(\tau_{H},\tau_{f}).$ (24)
$\tau_{C}$ and $\tau_{s}$ are the battery’s unitless energy and spin initial temperatures, where $\tau_{C}$ can be related to the absolute temperature $T_{C}$ by $\tau_{C}=T_{C}k_{B}/h\nu$ (Appendix E.3), and $\tau_{H}>\tau_{C},\tau_{s}$. $\tau_{f}$ is the unitless final temperature once equilibrium is reached, and is common across the conserved quantity reservoirs as required by the zero’th law (2). The final temperature is dependent on the entropic responses $C_{s}/\tau$ of the environment, and the energy/spin reservoirs of the battery, where the entropic response is the entropic equivalent of heat capacity, see Appendix E.
V Results
As we are interested in the theoretical potential of work extraction using the entropy battery, then we should use particle statistics that maximize the battery’s entropy capacity. At face value, it would appear as though distinguishable particles would provide the highest entropy capacity according to eq.(2), however, as will be discussed in §VII, distinguishability implies non-interaction, making distinguishable particles unrealistic as our model (and real systems) require interactions between the canonical ensembles for entropy to transfer. The next best case is an ensemble of bosons. If both the environment and the battery have a large number of bosons, on the order of Avogadro’s number, then the waste responses $C_{s},C_{V}$ in the integrals of eq.(24) are approximately eq.(2), and the indefinite integrals $\int C_{s}/\tau d\tau$ evaluate to
@@ V Results
We use this expression to numerically solve for $\tau_{f}$ by minimising $|\Delta\mathbb{S}|$ in eq.(24), although $\tau_{f}$ in the thermodynamic limit (large $N$ and large initial/final temperatures) does approximately follow a geometric mean $\tau_{f}\approx\sqrt[3]{\tau_{H}\tau_{C}\tau_{s}}$, consistent with the final temperature derived by a trio of reservoirs of unequal temperatures after reaching equilibrium [51]. See Appendix G for the proof.
The numerically calculated final temperature is used in the analytic boson heat expression (2) to calculate the energetic heat transferred from the environment to the energy bath of the battery, and therefore the energetic work liberated. This represents the unitless work that is made available during a conventional heat engine cycle
$W_{\text{conventional}}=|Q_{H}|-|Q_{C}|.$
As discussed in the previous sections, the spin reservoir also absorbs entropy, mediated by the Raman heat engine. The Raman heat engine simultaneously liberates heat energy in the battery as energetic work at the cost of spin labor. Therefore, the total unitless energetic work of the entropy battery becomes
$W_{\text{battery}}=|Q_{H}|-|Q_{C}|+|\mathcal{Q}_{s}|,$ (26)
[…]
@@
$\eta_{\text{energy}}=\frac{W_{\text{battery}}}{|Q_{H}|}=1-\frac{|Q_{C}|-|\mathcal{Q}_{s}|}{|Q_{H}|},$ (27)
@@
with the results shown in Fig.3 for varying initial battery temperatures. Our results show that once $d_{s}\geq 2$, the energy efficiency surpasses the Carnot limit. At 5 spin states, all initial temperatures considered reach close to $100\%$ energy efficiency. This occurs despite the probability distributions in each degree of freedom being initially states of maximum entropy (generalized Gibbs distribution), without induced quantum coherence. Further, as the entropy transfer occurs adiabatically, the heat engines are all endoreversible. The energy efficiency with 0 spin states follows $1-\sqrt{\tau_{C}/\tau_{H}}$, making our model consistent with the expected efficiency one would obtain when running the entropy battery at maximum power [52, 53, 51]. See Appendix G for further details.
@@
Figure 3: Efficiency of energetic work extraction $\eta_{\text{energy}}$ for varying number of spin states per particle calculated using the entropy and heat equations (25) and (2). In these simulations we used $d_{\text{env}}=d_{\text{batt},E}=400$ without loss in accuracy. Also shown are 4 sets of initial temperatures for the ensembles energy and spin baths, with $\tau_{\text{batt}}=\tau_{C}=\tau_{s}$, while the environments initial temperature was set to $\tau_{H}=0.6$ for all cases. These unitless temperatures were chosen as it would correspond to a temperature range between room temperature, 300 K, and 600 K with a Debye cut off frequency of 2121 THz, which is a typical maximum frequency for metals [50]. These initial temperatures also correspond to the spin reservoir having a polarization fidelity of roughly 99.97%. 300-600 K is a typical temperature range for coal and nuclear power plants, and concentrated solar farms. As the number of spin states increases, the energy efficiency approaches 100%, with all initial conditions reaching approximately this limit at 5 states. This is well above the Carnot efficiency limit $\eta_{\text{Carnot}}=1-\tau_{C}/\tau_{H}$ which is shown by the red bars when $d=0$, with values between 16%-50% for the varying initial ensemble temperatures.
VI Discussion
@@
The Carnot limit sets the maximum efficiency achievable by a heat engine operating between two thermal energy reservoirs of different temperatures. As mentioned in §III, a common strategy in quantum heat engines to surpass this limit is to induce quantum coherence within the thermal bath, either via squeezed thermal states [31, 32, 33] or other induced coherences [34, 35, 36]. All of these devices operate below the generalized Carnot limit [45], therefore there is no real quantum advantage. In contrast, the entropy battery achieves super-Carnot efficiencies in large ensembles using only maximum entropy states, characterised by their unitless temperatures, demonstrating that this limit can be surpassed only when an energetically hot source is coupled to an ensemble with multiple conserved quantities via multiple unitary heat engines.
@@
$\mathbb{W}=|Q_{H}|-|Q_{C}|+|\mathcal{Q}_{s}|-|\mathcal{L}|,$
where $|\mathcal{L}|=|\mathcal{Q}_{s}|$. The generalized efficiency of the battery when considering all resource costs therefore returns to the Carnot efficiency limit $\eta=1-|Q_{C}|/|Q_{H}|$, which proves that the entropy battery does not gain a ‘free-lunch’. Eq.(27) is therefore the energetic efficiency of the battery. If the initial preparation of spin labor within the battery requires energy, and if this energy cost is equal to the energy saved during the operation of the entropy battery, then the entropy battery would not have an advantageous net energy efficiency. However, re-polarising spin reservoirs does not necessarily require an energy input, as naturally occurring sources of spin coherence can instead be used to re-polarize the reservoir. Consequently, the entropy battery now has an additional advantage: it extends the range of usable natural resources for energetic work extraction beyond energy alone.
The unitary operation that couples the external energy and internal spin degrees of freedom is also controllable. Combining this controllability with the assumed mutual independence of the spin degree of freedom, the state of ‘charge’ of the battery would remain approximately constant even in an energetically hot environment. This makes the entropy battery useful in applications that require such controllability, with it effectively performing the job of a conventional battery, all the while storing not energy, but information in the form of coherence when charged. With this coherence stored across multiple conserved quantities but within the same $N_{\text{res}}$ particles, then a battery of this kind could in principal extract more energetic work per particle than storable by conventional batteries.
Similarly, the entropy battery can function as a heat storage device. When the environment-battery system thermalizes, the final unitless temperature $\tau_{f}$ is lower than the initial unitless temperature of the hot environment ($\tau_{f}<\tau_{H}$), meaning that the entropy battery absorbs and stores heat. Because the stored heat is distributed between the other non-energetic conserved quantities, the entropy battery achieves a higher density of heat storage than a single cold thermal bath. Such a mechanism could find applications in systems requiring efficient heat removal, such as spacecraft operating in vacuum or high-performance computing.
@@
This article focuses on an ideal scenario in which the battery-environment systems thermalize adiabatically. We briefly note the role of open-system effects here. Irreversibility arising from an injection of entropy from an external source, or non-independent conserved quantities, such that the total entropy is non-extensive, $\mathbb{S}_{\text{battery}}<\mathbb{S}_{\text{energy}}+\mathbb{S}_{\text{spin}}$ both effect the effective entropy capacity of the battery. Both of these effects increases the final unitless temperature once general equilibrium is reached, and decreases the extractable work. Beyond improving isolation or enhancing the mutual independence of the conserved quantities, one could compensate for this loss by increasing the number of states, however that analysis is beyond the scope of this article.
@@
A particular field that may apply the concepts developed in this article is quantum error correction. The decoherence of a qubit over time can be described as a thermalization process with a thermal environment [54, 55], and as such, one can imagine that having access to another conserved quantity reservoir that is independent of the basis of measurement (the qubit) within a single physical system (the ion, superconductor etc.) would allow for the ‘re-direction’ of decoherence into the unused degree of freedom. If possible, this would slow the decoherence time of the measured qubit, potentially leading to less dependence on error correcting qubits. Actual implementation of such an error correcting technique is also beyond the scope of this article, however NMR algorithmic cooling [41, 40] bears some similarities in terms of using ancillary qubits (the spin states) for reversibly cooling the measurement qubit (the energy states).
VII Maximizing entropy capacity
@@
The analysis presented in this article focuses on a single operating cycle, after which the environment ensemble is left cooler than its initial state, $\tau_{f}<\tau_{H}$. Repeated operation of the battery, assuming the environment is reset to its original temperature after each cycle, would allow continuous extraction of energetic work until the battery reaches the same unitless temperature as the now effectively isothermal environment. These extra cycles increase the amount of work extractable from the hot thermal energy bath. Therefore, it is of interest to understand how one could maximize the battery’s entropy capacity. According to eqs.(2) and (2), one could increase the number of particles $N$ or states $d$, as either of these increases the number of microstates $\Omega$ and therefore the entropy capacity. Another way would be to increase the number of independent conserved quantities. The dependence of the entropy capacity on these variables can be understood directly by application of the maximum entropy principle [6].
@@
The probability distribution that maximizes entropy, under the constraints of probability normalization $\sum_{\Omega}P(\Omega)=1$, Shannon entropy (28), and the $M$ total conserved quantities
@@
where $\gamma_{m}=\tau_{m}^{-1}$ and $c_{m}$ are the Lagrange multiplier and normalization constant for the $m$th conserved quantity. In this article, all ensembles are treated as canonical, with no particle exchange and therefore no associated chemical potential. The ensemble normalization $\prod_{m}c_{m}=Z$ is,
@@
For the conserved quantity $A_{m}$, the normalized probability distribution is
@@
The microstates $\Omega_{m}$ of the $m$th conserved quantity are now independent of the other conserved quantities. Therefore, mutual independence over the conserved quantities maximizes the ensembles entropy capacity. To further increase entropy capacity, we are limited to increasing the number of microstates associated with the $M$ conserved quantities.
VII.1 Distinguishability
Figure 4: An example entropy distribution over spin temperature for the different particle statistics, distinguishable, bosonic and fermionic with 7 states. In (a) and (b) we show example distributions with $N=4$ and $N=6$ particles respectively, illustrating how distinguishable and bosonic maximum entropies increase with NN without bound, while the fermionic maximum entropy decreases as saturation is approached $N=d$ where $\mathbb{S}_{\text{Fermi}}=0$ for all temperature. Maximum entropy, or the systems entropy capacity is given by the entropy at infinite temperature $\tau=\infty$, which we have labeled for the different ensemble statistics.